# Quantum stochastic thermodynamics in the mesoscopic-leads formulation

### v1

@@ I Introduction
Only a handful of methods exist that can address fluctuations of thermodynamic quantities in this far-from-equilibrium, strongly coupled setting. Non-equilibrium Green functions (NEGF) and scattering theory Blanter and Büttiker (2000); Esposito et al. (2009); Agarwalla et al. (2012); Esposito et al. (2015) are generally limited to slowly driven or weakly interacting systems Moskalets and Büttiker (2004); Moskalets (2014); Potanina et al. (2021), while path integral methods Aurell (2018); Funo and Quan (2018a, b) have been applied to numerically compute heat fluctuations in small, nonintegrable open quantum systems Kilgour et al. (2019); Popovic et al. (2021) but scaling this to larger systems, e.g. spin chains Fux et al. (2023), remains an open challenge. Another approach to strong-coupling thermodynamics exploits a Markovian embedding Woods et al. (2014) such as the reaction-coordinate method Iles-Smith et al. (2014); Strasberg et al. (2016); Newman et al. (2017); Anto-Sztrikacs and Segal (2021); Diba et al. (2023), where a non-Markovian open quantum system is incorporated into a larger Markovian one, leading to equivalent dynamics for the original open system. A powerful many-body extension of this approach is the mesoscopic-leads formulation of quantum transport Subotnik et al. (2009); Lacerda et al. (2023a, b); Gruss et al. (2016); Guimarães et al. (2016); Elenewski et al. (2017); Uzdin et al. (2018); Reichental et al. (2018); Chen et al. (2019); Brenes et al. (2020); Dzhioev and Kosov (2011); Ajisaka et al. (2012); Ajisaka and Barra (2013); Zelovich et al. (2014); Chen et al. (2014); Oz et al. (2020); Schwarz et al. (2016, 2018); Lotem et al. (2020); Elenewski et al. (2021); Wójtowicz et al. (2021), where macroscopic reservoirs are approximated by a finite number of damped modes in the same spirit as the pseudomode approach to open quantum systems Imamoglu (1994); Garraway (1997a, b); Tamascelli et al. (2018); Lambert et al. (2019); Mascherpa et al. (2020). The mesoscopic-leads formalism has recently been adapted to compute fluctuating charge currents in driven systems Brenes et al. (2023), but this method is not applicable to heat currents. Importantly, moreover, all the aforementioned approaches are tailored to compute the characteristic function, from which features of the probability distribution (e.g. of charge, heat etc.) can only be extracted by numerical differentiation or Fourier transform, with their associated errors and limitations.
@@
In this work, we introduce a method to directly sample the probability distributions of charge, heat, and entropy production in driven, strongly coupled open quantum systems far from equilibrium. Our approach exploits a mesoscopic-leads setup where the dynamics of the extended system is described by a Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation (ME) Lindblad (1976); Gorini et al. (1976). Monitoring the individual exchanges of energy quanta with the reservoirs results in a stochastic unravelling of the dynamics into quantum-jump trajectories. Each trajectory represents the state of the extended system conditioned on the measurement record Wiseman and Milburn (2009), and sampling such trajectories recovers the full counting statistics of currents and other observables Landi et al. (2024). In particular, we show that trajectory sampling enables reconstruction of the full distributions of heat, work, and entropy production, with both temporal and energy resolution. We focus on non-interacting fermionic systems, although our methods could be combined with tensor-network methods Brenes et al. (2020); Lacerda et al. (2023a) in the future to address the stochastic thermodynamics of arbitrary interacting systems. In this context, we demonstrate thermodynamic consistency of this method at the stochastic level by verifying three fluctuation theorems for entropy production: specifically, for the total, uncertainty, and martingale entropy production Manzano and Zambrini (2022); Manzano et al. (2019). As an application, we then study the full heat statistics of information erasure in a driven quantum dot coupled strongly to a fermionic reservoir. Our approach enables us to access the full heat distribution for arbitrary driving speeds and coupling strengths, going beyond previous work on finite-time quantum information erasure Miller et al. (2020); Van Vu and Saito (2022); Rolandi and Perarnau-Llobet (2022).
We emphasise that, to our knowledge, no other existing method allows for direct stochastic sampling of thermodynamic quantities in non-Markovian settings. Having access to individual trajectories not only gives additional insight beyond analysis of the cumulants, e.g. when examining the physical processes underpinning rare fluctuations Miller et al. (2020), but also unlocks the toolkit of continuous measurement theory for investigating measurement and feedback in non-Markovian quantum many-body systems.
@@
The paper is structured as follows: in Sec. II the mesoscopic-leads formulation for fermions is introduced. Then, in Sec. II.1 the unconditional state dynamics is discussed, and further specified for non-interacting systems (Sec. II.2). In Sec. III the conditional dynamics is presented. In Sec. IV, we discuss the thermodynamically consistent inference of particle (Sec. IV.0.1) and energy current (Sec. IV.0.2) along single trajectories, and the energy’s splitting into heat and work contributions (Sec. IV.0.3). In Sec. V, we discuss the total entropy production along a trajectory, relying on a two-point measurement scheme (TPM) in the mesoscopic-leads formalism. We demonstrate the emergence of integral fluctuation theorems for both the total entropy production and the uncertainty and martingale entropy production separately, in Sec. V.4. Subsequently, we study the fluctuations of dissipated heat during finite-time information erasure in Sec. VI.
@@ II.1 State dynamics
We consider a fermionic central system $\mathrm{S}$ described by a set of $N_{\mathrm{S}}$ annihilation operators $\{f_{j}\}_{j=1}^{N_{\mathrm{S}}}$. The system, with Hamiltonian $H_{\mathrm{S}}(t)$, may be interacting and driven externally. The system is coupled locally to $N_{\mathrm{R}}$ thermal reservoirs, indexed by $\alpha$, with Hamiltonians (in natural units, $\hbar=1$, $k_{\mathrm{B}}=1$) $H_{\mathrm{R_{\alpha}}}=\sum_{n=1}^{\infty}\omega_{\alpha,n}b_{\alpha,n}^{\dagger}b_{\alpha,n}$, via the interaction $H_{\mathrm{SR}_{\alpha}}(t)=\sum_{n=1}^{\infty}\lambda_{\alpha,n}(t)f^{\dagger}_{p_{\alpha}}b_{\alpha,n}+\mathrm{h.c.}$, where ${p_{\alpha}}$ denotes the index of the system site that the reservoir $\alpha$ couples to. The corresponding reservoir spectral density is given by $\mathcal{J}_{\alpha}(\omega)=2\pi\sum_{n=1}^{\infty}|\lambda_{\alpha,n}|^{2}\delta(\omega-\omega_{\alpha,n})$.
In the mesoscopic-leads formulation each thermal reservoir $\alpha$, at temperature $T_{\alpha}$ and chemical potential $\mu_{\alpha}$, is modeled by a mesoscopic lead $\alpha$ with a finite number $L_{\alpha}$ of fermionic modes with annihilation operators $\{a_{\alpha,k}\}_{\alpha=1,k=1}^{N_{\mathrm{R}},L_{\alpha}}$ and self-energies $\epsilon_{\alpha,k}$, so that $H_{\mathrm{L}_{\alpha}}=\sum_{k=1}^{L_{\alpha}}\epsilon_{\alpha,k}a_{\alpha,k}^{\dagger}a_{\alpha,k}$. The coupling between the system and each reservoir is replaced with the respective system-lead interaction $H_{\mathrm{SL}_{\alpha}}(t)=\sum_{k=1}^{L_{\alpha}}\kappa_{\alpha,k}(t)f^{\dagger}_{p_{\alpha}}a_{\alpha,k}+\mathrm{h.c.}$
Crucially, the residual reservoirs of each lead mode have a flat spectral density Garraway (1997a, b) and effective damping rate $\gamma_{\alpha,k}$ (typically one sets $\gamma_{\alpha,k}=\epsilon_{\alpha,k+1}-\epsilon_{\alpha,k}$russ et al. (2016); Elenewski et al. (2017); Gruss et al. (2017); Wójtowicz et al. (2020); Elenewski et al. (2021); Wójtowicz et al. (2021). Then, the coupling rate between the system site and modes in lead $\alpha$ is given by $\kappa_{\alpha,k}=\sqrt{\mathcal{J}_{\alpha}(\omega){\gamma_{\alpha,k}}/2\pi}$. Thus, each mode in the lead undergoes damping through interaction with a local Markovian reservoir, as illustrated schematically in Fig. 1.
For sufficiently large $L_{\alpha}$, $\gamma_{\alpha,k}$ becomes small, so that the residual reservoirs $(\mathrm{RR})$ may be traced out. The remaining extended system, described by the reduced density matrix $\rho=\mathrm{Tr}_{\mathrm{RR}}\left[\rho_{\mathrm{full}}\right]$, is composed of the system modes as well as the lead modes. Its dynamics is generated by the Liouvillian superoperator $\mathcal{L}$, so that $\frac{\mathrm{d}\rho}{\mathrm{d}t}=\mathcal{L}\left[\rho\right]$, in GKSL form
@@ II.2 State dynamics in non-interacting fermionic systems
with annihilation operators describing the extended system $\{c_{i}\}_{i=1}^{N_{\mathrm{S}}+L}$ and $L=\sum_{\alpha=1}^{N_{\mathrm{R}}}L_{\alpha}$. Many quantities of interest, such as thermodynamic currents in the extended system, may then be described solely in terms of the covariance matrix $C$, with size $\left((N_{\mathrm{S}}+L)\times(N_{\mathrm{S}}+L)\right)$ and matrix elements $C_{ij}=\langle c_{j}^{\dagger}c_{i}\rangle_{\rho}$, rather than the full state $\rho$, with size $(2^{N_{\mathrm{S}}+L}\times 2^{N_{\mathrm{S}}+L})$. Switching to the Heisenberg picture, the unconditional dynamics of the covariance matrix $C$ is governed by the ME
@@ III Quantum-jump trajectories in non-interacting fermionic systems
$\begin{split}\mathcal{L}\left[\rho\right]&=\left(\mathcal{L}_{0}+\mathcal{L}_{1}\right)\left[\rho\right],\\ \mathcal{L}_{0}\left[\rho\right]&=-i\left[H,\rho\right]-\frac{1}{2}\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{L_{\alpha}}\sum_{\sigma\in\{+,-\}}\{{L_{\alpha,k}^{\sigma}}^{\dagger}L_{\alpha,k}^{\sigma},\rho\},\\ \mathcal{L}_{1}\left[\rho\right]&=\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{L_{\alpha}}\sum_{\sigma\in\{+,-\}}L_{\alpha,k}^{\sigma}\rho{L_{\alpha,k}^{\sigma}}^{\dagger}.\end{split}$ (7)
Then the solution to Eq. (1) with initial state $\rho(t=0)=\rho_{0}$, generally given by $\rho(t)=e^{\mathcal{L}t}\rho_{0}$, may be expanded in a Dyson series as
@@ III.1 Gaussian initial states
Since now the state $\rho(t)$ is fully determined by its its covariance matrix $C(t)$, all formulas can only depend on matrices that are of size $\left((N_{\mathrm{S}}+L)\times(N_{\mathrm{S}}+L)\right)$, and thus they can be used efficiently, even for systems involving many modes. Furthermore, we emphasize that the subsequent findings are applicable to the stochastic jump dynamics of any fermionic non-interacting system subjected to thermal boundary driving, provided that the initial state is Gaussian.
@@ III.2 Conditional evolution
where $L^{+}_{k}=\sqrt{\gamma_{k}f_{k}}c_{k}^{\dagger}$ and $L^{-}_{k}=\sqrt{\gamma_{k}(1-f_{k})}c_{k}$ are the jump operators associated with absorption/emission events respectively and $\mathbb{1}$ denotes the identity operator. The measurement record is denoted by $r_{(0,\tau)}=\{(t_{J},k,\sigma);0<t_{J}<\tau\}$, recording a jump at time $t_{J}$ in lead mode $k$, where the jump signals particle transfer either on ($\sigma=+$) or off $(\sigma=-)$ it. Note that we neglect the probability of multiple jumps at one time. Thus, at every time $t\in(0,\tau)$ our knowledge about the system’s state must be updated: either conditioned on no-jump being observed or conditioned on a jump just having been registered.
@@
The stochastic increments of the point processes $\mathrm{d}N^{\sigma}_{k}$ obey Wiseman and Milburn (2009)

### v2

@@ I Introduction
Only a handful of methods exist that can address fluctuations of thermodynamic quantities in this far-from-equilibrium, strongly coupled setting. Non-equilibrium Green functions (NEGF) and scattering theory Blanter and Büttiker (2000); Esposito et al. (2009); Agarwalla et al. (2012); Esposito et al. (2015) are generally limited to slowly driven or weakly interacting systems Moskalets and Büttiker (2004); Moskalets (2014); Potanina et al. (2021), while path integral methods Aurell (2018); Funo and Quan (2018a, b) have been applied to numerically compute heat fluctuations in small, nonintegrable open quantum systems Kilgour et al. (2019); Popovic et al. (2021) but scaling this to larger systems, e.g. spin chains Fux et al. (2023), remains an open challenge. Another approach to strong-coupling thermodynamics exploits a Markovian embedding Woods et al. (2014) such as the reaction-coordinate method Iles-Smith et al. (2014); Strasberg et al. (2016); Newman et al. (2017); Anto-Sztrikacs and Segal (2021); Diba et al. (2024), where a non-Markovian open quantum system is incorporated into a larger Markovian one, leading to equivalent dynamics for the original open system. A powerful many-body extension of this approach is the mesoscopic-leads formulation of quantum transport Subotnik et al. (2009); Lacerda et al. (2024, 2023a); Gruss et al. (2016); Guimarães et al. (2016); Elenewski et al. (2017); Uzdin et al. (2018); Reichental et al. (2018); Chen et al. (2019); Brenes et al. (2020); Dzhioev and Kosov (2011); Ajisaka et al. (2012); Ajisaka and Barra (2013); Zelovich et al. (2014); Chen et al. (2014); Oz et al. (2020); Schwarz et al. (2016, 2018); Lotem et al. (2020); Elenewski et al. (2021); Wójtowicz et al. (2021), where macroscopic reservoirs are approximated by a finite number of damped modes in the same spirit as the pseudomode approach to open quantum systems Imamoglu (1994); Garraway (1997a, b); Tamascelli et al. (2018); Lambert et al. (2019); Mascherpa et al. (2020). The mesoscopic-leads formalism has recently been adapted to compute fluctuating charge currents in driven systems Brenes et al. (2023), but this method is not applicable to heat currents. Importantly, moreover, all the aforementioned approaches are tailored to compute the characteristic function, from which features of the probability distribution (e.g. of charge, heat etc.) can only be extracted by numerical differentiation or Fourier transform, with their associated errors and limitations.
@@
In this work, we introduce a method to directly sample the probability distributions of charge, heat, and entropy production in driven, strongly coupled open quantum systems far from equilibrium. Our approach exploits a mesoscopic-leads setup where the dynamics of the extended system is described by a Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation (ME) Lindblad (1976); Gorini et al. (1976). Monitoring the individual exchanges of energy quanta with the reservoirs results in a stochastic unravelling of the dynamics into quantum-jump trajectories. Each trajectory represents the state of the extended system conditioned on the measurement record Wiseman and Milburn (2009), and sampling such trajectories recovers the full counting statistics of currents and other observables Landi et al. (2024). In particular, we show that trajectory sampling enables reconstruction of the full distributions of heat, work, and entropy production, with both temporal and energy resolution. We focus on non-interacting fermionic systems, although our methods could be combined with tensor-network methods Brenes et al. (2020); Lacerda et al. (2024) in the future to address the stochastic thermodynamics of arbitrary interacting systems. In this context, we demonstrate thermodynamic consistency of this method at the stochastic level by verifying three fluctuation theorems for entropy production: specifically, for the total, uncertainty, and martingale entropy production Manzano and Zambrini (2022); Manzano et al. (2019). As an application, we then study the full heat statistics of information erasure in a driven quantum dot coupled strongly to a fermionic reservoir. Our approach enables us to access the full heat distribution for arbitrary driving speeds and coupling strengths, going beyond previous work on finite-time quantum information erasure Miller et al. (2020); Van Vu and Saito (2022); Rolandi and Perarnau-Llobet (2022).
@@
We emphasise that, to our knowledge, no other existing method allows for direct stochastic sampling of thermodynamic quantities in non-Markovian settings. Having access to individual trajectories not only gives additional insight beyond analysis of the cumulants, e.g. when examining the physical processes underpinning rare fluctuations Miller et al. (2020), but also unlocks the toolkit of continuous measurement theory for investigating measurement and feedback in non-Markovian quantum many-body systems. The paper is structured as follows: in Sec. II the mesoscopic-leads formulation for fermions is introduced. Then, in Sec. II.1 the unconditional state dynamics is discussed, and further specified for non-interacting systems (Sec. II.2). In Sec. III the conditional dynamics is presented. In Sec. IV, we discuss the thermodynamically consistent inference of particle (Sec. IV.0.1) and energy current (Sec. IV.0.2) along single trajectories, and the energy’s splitting into heat and work contributions (Sec. IV.0.3). In Sec. V, we discuss the total entropy production along a trajectory, relying on a two-point measurement scheme (TPM) in the mesoscopic-leads formalism. We demonstrate the emergence of integral fluctuation theorems for both the total entropy production and the uncertainty and martingale entropy production separately, in Sec. V.4. Subsequently, we study the fluctuations of dissipated heat during finite-time information erasure in Sec. VI.
@@ II.1 State dynamics
We consider a fermionic central system $\mathrm{S}$ described by a set of $N_{\mathrm{S}}$ annihilation operators $\{d_{j}\}_{j=1}^{N_{\mathrm{S}}}$. The system, with Hamiltonian $H_{\mathrm{S}}(t)$, may be interacting and driven externally. The system is coupled locally to $N_{\mathrm{R}}$ thermal reservoirs, indexed by $\alpha$, with Hamiltonians (in natural units, $\hbar=1$, $k_{\mathrm{B}}=1$) $H_{\mathrm{R_{\alpha}}}=\sum_{n=1}^{\infty}\omega_{\alpha,n}b_{\alpha,n}^{\dagger}b_{\alpha,n}$, via the interaction $H_{\mathrm{SR}_{\alpha}}(t)=\sum_{n=1}^{\infty}\lambda_{\alpha,n}(t)d^{\dagger}_{p_{\alpha}}b_{\alpha,n}+\mathrm{h.c.}$, where $p_{\alpha}$} denotes the index of the system site that the reservoir $\alpha$ couples to. The corresponding reservoir spectral density is given by $\mathcal{J}_{\alpha}(\omega)=2\pi\sum_{n=1}^{\infty}|\lambda_{\alpha,n}|^{2}\delta(\omega-\omega_{\alpha,n})$.
In the mesoscopic-leads formulation each thermal reservoir $\alpha$, at temperature $T_{\alpha}$ and chemical potential $\mu_{\alpha}$, is modeled by a mesoscopic lead $\alpha$ with a finite number $l_{\alpha}$ of fermionic modes, each with a residual reservoir acting on it. The lead modes are described with annihilation operators $\{a_{\alpha,k}\}_{\alpha=1,k=1}^{N_{\mathrm{R}},l_{\alpha}}$ and have self-energies $\epsilon_{\alpha,k}$, so that $H_{\mathrm{L}_{\alpha}}=\sum_{k=1}^{l_{\alpha}}\epsilon_{\alpha,k}a_{\alpha,k}^{\dagger}a_{\alpha,k}$. The coupling between the system and each reservoir is replaced with the respective system-lead interaction $H_{\mathrm{SL}_{\alpha}}(t)=\sum_{k=1}^{l_{\alpha}}\kappa_{\alpha,k}(t)d^{\dagger}_{p_{\alpha}}a_{\alpha,k}+\mathrm{h.c.}$ Crucially, the residual reservoirs of each lead mode have a flat spectral density Garraway (1997a, b) and effective damping rate $\gamma_{\alpha,k}$ (typically one sets $\gamma_{\alpha,k}=\epsilon_{\alpha,k+1}-\epsilon_{\alpha,k}$Gruss et al. (2016); Elenewski et al. (2017); Gruss et al. (2017); Wójtowicz et al. (2020); Elenewski et al. (2021); Wójtowicz et al. (2021). Then, the coupling rate between the system site and modes in lead $\alpha$ is given by $\kappa_{\alpha,k}=\sqrt{\mathcal{J}_{\alpha}(\omega){\gamma_{\alpha,k}}/2\pi}$. Thus, each mode in the lead undergoes damping through interaction with a local Markovian reservoir, as illustrated schematically in Fig. 1.
For sufficiently large $l_{\alpha}$, $\gamma_{\alpha,k}$ becomes small, so that the residual reservoirs $(\mathrm{RR})$ may be traced out. The remaining extended system, described by the reduced density matrix $\rho$, is composed of the system modes as well as the lead modes. Its dynamics is generated by the Liouvillian superoperator $\mathcal{L}$, so that $\frac{\mathrm{d}\rho}{\mathrm{d}t}=\mathcal{L}\left[\rho\right]$, in GKSL form
@@ II.2 State dynamics in non-interacting fermionic systems
with annihilation operators describing the extended system $\{c_{i}\}_{i=1}^{N_{\mathrm{S}}+L}$ and $L=\sum_{\alpha=1}^{N_{\mathrm{R}}}l_{\alpha}$, so that $\{f_{1},\dots,f_{N_{\mathrm{S}}},a_{1,1},\dots,a_{1,l_{1}},\dots a_{N_{\mathrm{R}},1},\dots,a_{N_{\mathrm{R}},l_{N_{\mathrm{R}}}}\}\to\{c_{1},\dots,c_{N_{\mathrm{S}}+L}\}$. Many quantities of interest, such as thermodynamic currents in the extended system, may then be described solely in terms of the covariance matrix $C$, with size $\left((N_{\mathrm{S}}+L)\times(N_{\mathrm{S}}+L)\right)$ and matrix elements $C_{ij}=\langle c_{j}^{\dagger}c_{i}\rangle_{\rho}$, rather than the full state $\rho$, with size $(2^{N_{\mathrm{S}}+L}\times 2^{N_{\mathrm{S}}+L})$. Switching to the Heisenberg picture, the unconditional dynamics of the covariance matrix $C$ is governed by the ME
@@ III Quantum-jump trajectories in non-interacting fermionic systems
$\begin{split}\mathcal{L}\left[\rho\right]&=\left(\mathcal{L}_{0}+\mathcal{L}_{1}\right)\left[\rho\right],\\ \mathcal{L}_{0}\left[\rho\right]&=-i\left[H,\rho\right]-\frac{1}{2}\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{l_{\alpha}}\sum_{\sigma\in\{+,-\}}\left\{{L_{\alpha,k}^{\sigma}}^{\dagger}L_{\alpha,k}^{\sigma},\rho\right\},\\ \mathcal{L}_{1}\left[\rho\right]&=\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{l_{\alpha}}\sum_{\sigma\in\{+,-\}}L_{\alpha,k}^{\sigma}\rho{L_{\alpha,k}^{\sigma}}^{\dagger}.\end{split}$ (7)
Then the solution to Eq. (1) with initial state $\rho(t=0)=\rho_{0}$, generally given by $\rho(t)=e^{\mathcal{L}t}\rho_{0}$ for a time-independent $\mathcal{L}$, may be expanded in a Dyson series as
@@ III.1 Gaussian initial states
Since now the state $\rho(t)$ is fully determined by its covariance matrix $C(t)$, all formulas can only depend on matrices that are of size $\left((N_{\mathrm{S}}+L)\times(N_{\mathrm{S}}+L)\right)$, and thus they can be used efficiently, even for systems involving many modes. Furthermore, we emphasize that the subsequent findings are applicable to the stochastic jump dynamics of any fermionic non-interacting system subjected to thermal boundary driving, provided that the initial state is Gaussian.
@@ III.2 Conditional evolution
where $L^{+}_{k}=\sqrt{\gamma_{k}f_{k}}c_{k}^{\dagger}$ and $L^{-}_{k}=\sqrt{\gamma_{k}(1-f_{k})}c_{k}$ are the jump operators associated with absorption/emission events respectively and $\mathbb{1}$ denotes the identity operator. Note that we have combined the lead index $\alpha$ with the index specifying the mode within the corresponding lead into a single index $k$, which runs over all modes across all leads. The measurement record is denoted by $r_{(0,\tau)}=\{(t_{J},k,\sigma);0<t_{J}<\tau\}$, recording a jump at time $t_{J}$ in lead mode $k$, where the jump signals particle transfer either on ($\sigma=+$) or off $(\sigma=-)$ it. Note that we neglect the probability of multiple jumps at one time. Thus, at every time $t\in(0,\tau)$ our knowledge about the system’s state must be updated: either conditioned on no-jump being observed or conditioned on a jump just having been registered.
@@
The superoperators $\mathcal{G}$ and $\mathcal{H}$ describe two types of measurement backaction: $\mathcal{G}$ corresponds to instantaneous state updates that occur when a jump is registered ($\mathrm{d}N_{k}^{\sigma}=1$), while $\mathcal{H}$ represents the continuous, smooth effect between jumps ($\mathrm{d}N_{k}^{\sigma}=0$), since the absence of jump detection still provides information about the system. The stochastic increments of the point processes $\mathrm{d}N^{\sigma}_{k}$ obey Wiseman and Milburn (2009)